### v1

# A Proof of Gross’ Conjecture on $2$-Automorphic $2$-Groups

Abstract.
Building upon previous results, a classification is given of finite $p$-groups whose elements of order $p$ are all fused. This particularly confirms a conjecture of Gross proposed in 1976 on $2$-automorphic $2$-groups, which are $2$-groups with involutions forming a single fusion classes. As a consequence, two open problems regarding AT-groups and FIF-groups are solved.
Key words and phrases:
Suzuki $2$-groups; $2$-automorphic $2$-groups; exterior squre; dual module
1. Introduction
It is well-known that finite $p$-groups with a single subgroup of order $p$ are cyclic or a generalized quaternion group. The class of finite $p$-groups with a single fusion class of elements of order $p$ has been extensively studied. (Recall that a fusion class of a group $N$ is an orbit of ${\mathrm{Aut}}(N)$ on the elements.) The study was initiated by G. Higman 1963 when he classified finite non-abelian $2$-groups (with more than one involution) that has a cyclic group of automorphisms transitive on the involutions, called Suzuki $2$-groups.
For odd primes $p$, Shult [22] showed that homocyclic $p$-groups are the only $p$-groups of which elements of order $p$ form a single fusion class. Shaw [21] showed that a finite non-abelian $2$-group (with more than one involution) which has a solvable group of automorphisms transitive on the involutions is a Suzuki $2$-group. Higman’s classification associated with Shaw’s work forms Sections 5-7 of Huppert’s book [12, Chapter VIII].
In 1976, Gross [10] studied finite $2$-groups with a single fusion class of involutions, called $2$-automorphic $2$-groups. He further proposed a conjecture that (roughly speaking) $2$-automorphic $2$-groups are Suzuki $2$-groups. This conjecture has been partially solved by Bryukhanova [3] and Wilkens [24], and is recorded as Problem 7.31 in the Kourovka notebook [18].
Building upon previous results, the main theorem of this paper obtains a classification of finite $p$-groups of which elements of order $p$ form a single fusion class, completing the long-term classification problem and confirming Gross’ conjecture.
Theorem 1.1.
Let $N$ be a finite $p$-group such that ${\mathrm{Aut}}(N)$ is transitive on the set of all elements of order $p$. Then one of the following statements holds:
  1. (i)
    $N$ is a homocyclic $p$-group;
  2. (ii)
    $N$ is a generalized quaternion $2$-group;
  3. (iii)
    $N$ is a Suzuki $2$-group, and ${\mathrm{Aut}}(N)$ is solvable with a cyclic subgroup acting transitively on the set of involutions.
The proof of Theorem 1.1 for $2$-groups depends on certain representations of finite transitive linear groups over ${\mathbb{F}}_{2}$, stated below, which is independently interesting.
Theorem 1.2.
Let $V={\mathbb{F}}_{2}^{n}$, $G\leqslant{\mathrm{GL}}(V)={\mathrm{GL}}_{n}(2)$ be non-solvable, and $M=\Lambda^{2}_{{\mathbb{F}}_{2}}(V)/W$ for some ${\mathbb{F}}_{2}G$-submodule $W$ of $\Lambda^{2}_{{\mathbb{F}}_{2}}(V)$. Assume that $\dim M=\dim V$, the natural action of $G$ on $V$ is transitive on the non-zero vectors of $V$, and the induced action of $G$ on $M$ is transitive on the non-zero vectors of $M$. Then $n=3m$, $G^{(\infty)}={\mathrm{SL}}_{3}(2^{m})<{\mathrm{GL}}_{3m}(2)$, and $M$ is the dual ${\mathbb{F}}_{2}G^{(\infty)}$-module of $V$.
Solving Gross’ conjecture is also an important part for the study of some other classes of groups. A group $N$ is called an AT-group if ${\mathrm{Aut}}(N)$ is transitive on the elements of the same order. Zhang [25] presents a description of finite AT-groups. However, the problem of determining AT-$2$-groups has been left open, as commented that “If Gross’ conjecture on $2$-automorphic $2$-groups is true, …… We do not know if Suzuki $2$-groups are AT-groups” before Lemma 2.2 of [25]. A group is called an FIF-group if any two elements of the same order are fused or inverse-fused. Praeger and the first-named author [16] give a description of finite FIF-groups, with the problem of classifying FIF-$2$-groups left open, see [16, Problem 1.5]. As a corollary of Theorem 1.1, these problems are solved.
Corollary 1.3.
Let $N$ be a non-abelian finite $2$-group with more than one involution. Then the following statements are equivalent:
  1. (i)
    cyclic subgroups of the same order are conjugate under ${\mathrm{Aut}}(N)$;
  2. (ii)
    any two elements of $N$ of the same order are conjugate or inverse-conjugate under ${\mathrm{Aut}}(N)$;
  3. (iii)
    elements of $N$ of the same order are conjugate under ${\mathrm{Aut}}(N)$;
  4. (iv)
    $N=A_{2}(n,\theta)$, $B_{2}(n)$, or $P(\epsilon)$; see Definitions 4.1, 4.3 and 4.5.
We remark that the automorphism groups of the groups $A_{2}(n,\theta)$, $B_{2}(n)$ and $P(\epsilon)$ are determined in Lemmas 4.2, 4.4 and 4.7.
Acknowledgments
The authors acknowledge the support of NNSFC grant no. 11931005.
2. Transitive linear groups and their exterior squares
All fields and groups are finite in this section. For a field ${\mathrm{F}}$ and a group $G$, an ${\mathrm{F}}G$-reprensentation is a group homomorphism $\rho$ from $G$ to ${\mathrm{GL}}(V)$ for some ${\mathrm{F}}$-space $V$. Every ${\mathrm{F}}G$-reprensentation $\rho$ induces an ${\mathrm{F}}G$-module structure on $V$ by $(xg)\cdot v=x\cdot v^{\rho(g)}$ for $x\in{\mathrm{F}}$ and $g\in G$. If ${\mathrm{E}}$ is an extension field of ${\mathrm{F}}$, then $V^{\mathrm{E}}:=V\otimes{\mathrm{E}}$ is naturally an ${\mathrm{E}}G$-module. Some elementary properties of the operator $\cdot\otimes{\mathrm{E}}$ can be found in [12, Section VII.1], [1, Chapter 9] and [13, Chapter 9]. For an ${\mathrm{F}}G$-module $V$ with representation $\rho$, we use the following two notations:
  • denote by $V^{\sigma}$ the ${\mathrm{F}}G$-module induced by $\sigma\circ\rho$ for any field automorphism $\sigma$ of ${\mathrm{F}}$;
  • denote by $V_{\mathrm{K}}$ the ${\mathrm{K}}G$-module of $V$ for subfield ${\mathrm{K}}$ of ${\mathrm{F}}$.
Some important propositions are recorded blow.
Lemma 2.1.
Suppose that ${\mathrm{K}}\subset{\mathrm{F}}\subset{\mathrm{E}}$ are field extensions, and $V,W$ are ${\mathrm{F}}G$-modules. Then the following statements hold:
  1. (i)
    [12, Chapter VII, 1.16(a)] $(V_{\mathrm{K}})^{\mathrm{F}}\cong\sum_{\sigma\in{\mathrm{Gal}}({\mathrm{F}}/{\mathrm{K}})}V^{\sigma}$;
  2. (ii)
    [1, 25.7(2)] if there exists an ${\mathrm{E}}G$-submodule $U$ of $V^{\mathrm{E}}$ such that $U^{\sigma}=U$ for any $\sigma\in{\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$, then $U=(V_{0})^{\mathrm{E}}$ for some ${\mathrm{F}}G$-submodule $V_{0}$ of $V$;
  3. (iii)
    [Deuring-Noether, see [12, Chapter VII, 1.22]] $V^{\mathrm{E}}\cong W^{\mathrm{E}}$ if and only if $V\cong W$.
The exterior square $\Lambda^{2}_{{\mathrm{F}}}(V)$ of an ${\mathrm{F}}G$-module $V$ is $(V\otimes V)/S$ with $S=\langle v\otimes v:v\in V\rangle$. Note that if $V=V_{1}\oplus\cdots\oplus V_{n}$, we have
$\Lambda^{2}_{{\mathrm{F}}}(V)\cong\bigoplus_{i=1}^{n}\Lambda_{{\mathrm{F}}}^{2}(V_{i})\oplus\bigoplus_{1\leqslant i<j\leqslant n}V_{i}\otimes V_{j}.$
Then we can prove the following lemma for decomposing exterior squares.
Lemma 2.2.
Suppose that $V$ is an ${\mathrm{F}}G$-module and ${\mathrm{E}}$ is an extension field of ${\mathrm{F}}$. If $U$ is an ${\mathrm{E}}G$-module such that $U_{{\mathrm{F}}}\cong V$, then the following statements hold:
  1. (i)
    $V^{\mathrm{E}}\cong\bigoplus_{i=1}^{f}U^{\phi^{i}}$ where $\phi$ is a generator of ${\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$ and $f=[{\mathrm{E}}:{\mathrm{F}}]$;
  2. (ii)
    $\Lambda^{2}_{\mathrm{F}}(V)=A\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}$ where
    1. (1)
      $A\cong\Lambda_{\mathrm{E}}^{2}(U)_{\mathrm{F}}$;
    2. (2)
      $B_{j}\cong\bigl{(}U\otimes U^{\phi^{j}}\bigr{)}_{\mathrm{F}}$, when $21\leqslant j<f/$ 2;
    3. (3)
      $2\cdot B_{f/2}\cong\bigl{(}U\otimes U^{\phi^{f/2}}\bigr{)}_{\mathrm{F}}$, when $f$ is even.
Proof.
Part (i) is directly deduced by Lemma 2.1(i).
Note that $\Lambda^{2}_{\mathrm{F}}(V)=(V\otimes V)/S$ with $S=\langle v\otimes v:v\in V\rangle$, then $\Lambda^{2}_{\mathrm{F}}(V)^{\mathrm{E}}\cong(V\otimes V)^{\mathrm{E}}/S^{\mathrm{E}}\cong\Lambda_{\mathrm{E}}^{2}(V^{\mathrm{E}})$. Thus, we obtain that
$\Lambda^{2}_{\mathrm{F}}(V)^{\mathrm{E}}$ $\cong\Lambda_{\mathrm{E}}^{2}(V^{\mathrm{E}})\cong\Lambda_{\mathrm{E}}^{2}(U\oplus U^{\phi}\oplus\cdots\oplus U^{\phi^{f-1}})$
$\cong\bigoplus_{i=1}^{f}\Lambda_{\mathrm{E}}^{2}(U)^{\phi^{i}}\oplus\bigoplus_{1\leqslant i<j\leqslant f}U^{\phi^{i}}\otimes U^{\phi^{j}}.$
We use the notations:
  • $\widetilde{A}\cong\bigoplus_{i=1}^{f}\Lambda_{\mathrm{E}}^{2}(U)^{\phi^{i}}$;
  • $\widetilde{B}_{j}\cong\bigoplus_{i=1}^{f}\bigl{(}U\otimes U^{\phi^{j}}\bigr{)}^{\phi^{i}}$ for $1\leqslant j<\frac{f}{2}$; and
  • $\widetilde{B}_{f/2}=\bigoplus_{i=1}^{f/2}\bigl{(}U\otimes U^{\phi^{f/2}}\bigr{)}^{\phi^{i}}$ when $f$ is even.
Hence, $\Lambda^{2}_{\mathrm{F}}(V)^{\mathrm{E}}\cong\widetilde{A}\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}\widetilde{B}_{j}$ by rearranging the summands. Note that $\phi$ fixes $\widetilde{A}$ and $\widetilde{B}_{j}$ for each $1\leqslant j\leqslant\frac{f}{2}$, then there exist ${\mathrm{F}}G$-submodules $A$ and $B_{j}$’s of $\Lambda^{2}_{\mathrm{F}}(V)$ such that $A^{\mathrm{E}}\cong\widetilde{A}$ and $B_{j}^{\mathrm{E}}\cong\widetilde{B}_{j}$ by Lemma 2.1(ii). Thus, we have $\Lambda^{2}_{\mathrm{F}}(V)=A\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}$ by Lemma 2.1(iii).
Lemma 2.1(i) deduces that
$(\Lambda_{\mathrm{E}}^{2}(U)_{\mathrm{F}})^{\mathrm{E}}\cong\bigoplus_{i=1}^{f}\Lambda_{\mathrm{E}}^{2}(U)^{\phi^{i}}=\widetilde{A}\cong A^{\mathrm{E}}.$
Thus, $A\cong\Lambda_{\mathrm{E}}^{2}(U)_{\mathrm{F}}$ by Lemma 2.1(iii), as in (1) of part (ii). Similarly, we obtain that $B_{j}\cong\bigl{(}U\otimes U^{\phi^{j}}\bigr{)}_{\mathrm{F}}$, when $21\leqslant j<f/$ 2, as in (2) of part (ii). When $j$ is even, we have
$\left(\bigl{(}U\otimes U^{\phi^{f/2}}\bigr{)}_{\mathrm{F}}\right)^{\mathrm{E}}\cong\bigoplus_{i=1}^{f}\bigl{(}U\otimes U^{\phi^{f/2}}\bigr{)}^{\phi^{i}}\cong 2\cdot\bigoplus_{i=1}^{f/2}\bigl{(}U\otimes U^{\phi^{f/2}}\bigr{)}^{\phi^{i}}\cong 2\cdot\widetilde{B}_{f/2}.$
Hence, $2\cdot B_{f/2}\cong U\otimes U^{\phi^{f/2}}$ deduced by Lemma 2.1(iii), as in (3) of part (ii). ∎
Remark 2.3.
Let $V$ be an irreducible ${\mathrm{F}}G$-module. Then ${\mathrm{E}}:={\mathrm{End}}_{{\mathrm{F}}G}(V)$ is an extension field of ${\mathrm{F}}$ for $V$ by Schur’s Lemma and Wedderburn’s Little Theorem. Then $V$ has an ${\mathrm{E}}G$-module structure defined by $\alpha g:v\mapsto v^{\alpha g}=v^{g\alpha}$ for $\alpha\in{\mathrm{E}}$ and $g\in G$. Let $U$ be the ${\mathrm{E}}G$-module of $V$ defined as above. Then $U_{\mathrm{F}}\cong V$. Hence, the decomposition of exterior square of an irreducible ${\mathrm{F}}G$-modules can be well-treated by extending the field to ${\mathrm{End}}_{{\mathrm{F}}G}(V)$.
Let $V={\mathbb{F}}_{2}^{n}$ be a vector space, and let $G\leqslant{\mathrm{GL}}(V)\cong{\mathrm{GL}}_{n}(2)$ be transitive on $V\setminus\{0\}$. These groups $G$ are known by the classification of the finite $2$-transitive permutation groups, and we list them in the next theorem, see [4, Table 7.3].
Theorem 2.4.
Suppose that $G\leqslant{\mathrm{GL}}_{n}(2)$ acts transitively on non-zero vectors of $V={\mathbb{F}}_{2}^{n}$. Then one of the following statements holds:
  1. (i)
    $G\leqslant{\Gamma\mathrm{L}}_{1}(2^{n})$ is solvable;
  2. (ii)
    $(G,n)=({\mathrm{A}}_{6},4)$, $({\mathrm{Sp}}_{4}(2),4)$, $({\mathrm{A}}_{7},4)$, $({\mathrm{PSU}}_{3}(3),6)$ or $({\mathrm{G}}_{2}(2),6)$;
  3. (iii)
    $G^{(\infty)}$ is transitive on non-zero vectors and is one of the following groups:
    1. (1)
      ${\mathrm{SL}}_{m}(2^{f})$ with $m\geqslant 2$ and $(m,f)\neq(2,1)$;
    2. (2)
      ${\mathrm{Sp}}_{2m}(2^{f})$ with $m\geqslant 2$ and $(m,f)\neq(2,1)$; or
    3. (3)
      ${\mathrm{G}}_{2}(2^{f})$ with $f\geqslant 2$.
The following lemma can be directly deduced by Steinberg’s Twisted Tensor Product Theorem [23] (also see [14, Theorem 5.4.1]).
Lemma 2.5.
Suppose that $G\cong{\mathrm{Sp}}_{2m}(p^{f})$, ${\mathrm{SL}}_{m}(p^{f})$ or ${\mathrm{G}}_{2}(p^{f})$ acting naturally on ${\mathbb{F}}_{p^{f}}$-space $U$ for prime $p$. Let $\phi:x\mapsto x^{p}$ be the Frobenius automorphism of ${\mathbb{F}}_{p^{f}}$. For $11\leqslant i,j,k,\ell\leqslant f-$ 1, the following statements hold:
  1. (i)
    $U^{\phi^{i}}\otimes U^{\phi^{j}}$ is an irreducible ${\mathbb{F}}_{p^{f}}G$-module for $i\neq j$;
  2. (ii)
    for $i\neq j$ and $k\neq\ell$, $U^{\phi^{i}}\otimes U^{\phi^{j}}\cong U^{\phi^{k}}\otimes U^{\phi^{\ell}}$ if and only if $\{i,j\}=\{k,\ell\}$.
We say an ${\mathrm{E}}G$-module $U$ can be written over a subfield ${\mathrm{F}}\leqslant{\mathrm{E}}$ if the matrix of each $g\in G$, with respect to some basis, has entries in ${\mathrm{F}}$. Hence, $U$ can be written over ${\mathrm{F}}$ if and only if $U^{\sigma}\cong U$ for any $\sigma\in{\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$. In addition, $U_{\mathrm{K}}$ is irreducible for any subfield ${\mathrm{F}}$ of ${\mathrm{E}}$ if and only if $U$ cannot be written over any subfield, see [1, 26.5]. Immediately we obtain the following lemma.
Lemma 2.6.
Suppose that $G\cong{\mathrm{Sp}}_{2m}(q)$, ${\mathrm{SL}}_{m}(q)$ or ${\mathrm{G}}_{2}(q)$ with $q=p^{f}$ for some prime $p$. Let $V$ be the natural ${\mathbb{F}}_{p}G$-module, and let $U$ be the natural ${\mathbb{F}}_{q}G$-module. Then
  1. (i)
    $V^{{\mathbb{F}}_{q}}\cong U\oplus U^{\phi}\oplus\cdots\oplus U^{\phi^{f-1}}$, where $\phi:x\mapsto x^{p}$ the Frobenius automorphism of ${\mathbb{F}}_{q}$;
  2. (ii)
    $(U\otimes U^{\phi^{j}})_{{\mathbb{F}}_{p}}$ is irreducible for $j<\frac{f}{2}$;
  3. (iii)
    when $f$ is even, $(U\otimes U^{\phi^{f/2}})_{{\mathbb{F}}_{p}}$ is a direct sum of two isomorphic copies of irreducible ${\mathbb{F}}_{p}G$-modules.
Proof.
For convenience, let ${\mathrm{E}}={\mathbb{F}}_{q}$, ${\mathrm{F}}={\mathbb{F}}_{p}$, and let ${\mathrm{L}}={\mathbb{F}}_{p^{f/2}}$ when $f$ is even.
It is clear that $U_{\mathrm{F}}\cong V$, and then part (i) holds by Lemma 2.2(i).
Suppose that $21\leqslant j<f/$ 2. Then we have $(U\otimes U^{\phi^{j}})^{\phi^{k}}\cong U\otimes U^{\phi^{i}}$ if and only if $f\mid k$ by lemma 2.5. Hence, $U\otimes U^{\phi^{j}}$ cannot be written over any subfield of ${\mathrm{E}}$, and thus $(U\otimes U^{\phi^{j}})_{\mathrm{F}}$ is irreducible, as in part (ii).
Assume from now that $f$ is even. Note that ${\mathrm{Gal}}({\mathrm{E}}/{\mathrm{L}})=\langle\phi^{f/2}\rangle$ and $U\otimes U^{\phi^{f/2}}$ is normalized by $\phi^{f/2}$. By Lemma 2.1(ii), there exists an ${\mathrm{L}}G$-module $W$ such that $W^{{\mathrm{E}}}\cong U\otimes U^{\phi^{f/2}}$. Lemma 2.1(i) deduces that
$((U\otimes U^{\phi^{f/2}})_{\mathrm{L}})^{{\mathrm{E}}}\cong(U\otimes U^{\phi^{f/2}})\oplus(U\otimes U^{\phi^{f/2}})^{\phi^{f/2}}\cong 2\cdot U\otimes U^{\phi^{f/2}}\cong 2\cdot W^{{\mathrm{E}}}.$
Hence, $(U\otimes U^{\phi^{f/2}})_{\mathrm{L}}=2\cdot W$ by Lemma 2.1(iii). In particular, we have
$(U\otimes U^{\phi^{f/2}})_{{\mathrm{F}}}=((U\otimes U^{\phi^{f/2}})_{{\mathrm{L}}})_{{\mathrm{F}}}=2\cdot W_{\mathrm{F}}.$
We only need to show that $W_{\mathrm{F}}$ is an irreducible ${\mathrm{F}}G$-module, that is, $W$ cannot be written over any subfield of ${\mathrm{L}}$.
Let $\psi$ be the image of $\phi$ in ${\mathrm{Gal}}({\mathrm{L}}/{\mathrm{F}})$. Then ${\mathrm{Gal}}({\mathrm{L}}/{\mathrm{F}})=\langle\psi\rangle$. Suppose that $W^{\psi^{k}}\cong W$ for some $k$. We have
$W^{\psi^{k}}\cong W$ $\Rightarrow(2\cdot W)^{\psi^{k}}\cong 2\cdot W\Rightarrow((U\otimes U^{\phi^{f/2}})_{\mathrm{L}})^{\psi^{k}}\cong(U\otimes U^{\phi^{f/2}})_{\mathrm{L}}$
$\Rightarrow((U\otimes U^{\phi^{f/2}})^{\phi^{k}})_{\mathrm{L}}\cong(U\otimes U^{\phi^{f/2}})_{\mathrm{L}}$
$\Rightarrow((U^{\phi^{k}}\otimes U^{\phi^{f/2+k}})_{\mathrm{L}})^{{\mathrm{E}}}\cong((U\otimes U^{\phi^{f/2}})_{\mathrm{L}})^{{\mathrm{E}}}$
$\Rightarrow 2\cdot U^{\phi^{k}}\otimes U^{\phi^{f/2+k}}\cong 2\cdot U\otimes U^{\phi^{f/2}}.$
Note that $U^{\phi^{k}}\otimes U^{\phi^{f/2+k}}\cong U\otimes U^{\phi^{f/2}}$ if and only if $k$ is divisible by $\frac{f}{2}$ by Lemma 2.5. Hence, we have $\psi^{k}=1$, and then $W$ cannot be written over any subfield of ${\mathrm{L}}$. It follows that $W_{{\mathrm{F}}}$ is an irreducible ${\mathrm{F}}G$-module, which proves part (iii). ∎
We record some classical results for exterior squares of natural modules for ${\mathrm{SL}}_{m}(q)$, ${\mathrm{Sp}}_{2m}(q)$ and ${\mathrm{G}}_{2}(q)$.
Lemma 2.7 ([17, Theorem 2.2]).
Let $U={\mathbb{F}}_{q}^{m}$, and let $G={\mathrm{SL}}_{m}(q)$. Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is an irreducible ${\mathbb{F}}_{q}G$-module which cannot be written over an subfield of ${\mathbb{F}}_{q}$.
In particular, if $m=3$, then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is isomorphic to the dual ${\mathbb{F}}_{q}G$-module of $V$.
Results of exterior square of the natural modules of symplectic groups can be found in [5, 15, 19], and the result of the exceptional group of Lie type ${\mathrm{G}}_{2}(q)<{\mathrm{Sp}}_{6}(q)$ (even $q$) can be deduced from [20, Theorem 8.14].
Lemma 2.8.
Suppose that $U={\mathbb{F}}_{q}^{2m}$ is a symplectic space with $q=2^{f}$, and $G={\mathrm{Sp}}_{2m}(q)$ or ${\mathrm{G}}_{2}(q)$($m=3$). Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ has a maximal ${\mathbb{F}}_{q}G$-submodule $T$ of codimension $1$. Moerover, $T$ is indecomposable and one the following statements holds:
  1. (i)
    if $p\nmid m$ then $T$ is irreducible and cannot be written over any subfield;
  2. (ii)
    if $p\mid m$, then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is indecomposable and $T$ has a unique maximal ${\mathbb{F}}_{q}G$-submodule $T_{0}$ which is trivial, and $T/T_{0}$ cannot be written over any subfield.
Now we are ready to prove Theorem 1.2.
Proof of Theorem 1.2.
Let $G\leqslant{\mathrm{GL}}(V)$ with $V={\mathbb{F}}_{2}^{n}$, and let $M=\Lambda^{2}_{{\mathbb{F}}_{2}}(V)/W$ for a submodule $W$. Assume that $G$ is transitive on $V\setminus\{0\}$. Then $G$ is listed in Theorem 2.4. If $G$ is one of the small groups listed in Theorem 2.4(ii), then computations in Magma [2] shows that $\dim M=1$. Thus, we only need to focus on the cases that $G^{(\infty)}\cong{\mathrm{SL}}_{m}(q)$ ($m\geqslant 2$ and $(m,f)\neq(2,1)$), ${\mathrm{Sp}}_{2m}(q)$ ($m\geqslant 2$ and $(m,q)\neq(2,2)$) or ${\mathrm{G}}_{2}(q)$ ($q\geqslant 4$) for $q=2^{f}$.
Let $K$ be the kernel of $G$ acting on $M$. Then $G/K$ is faithful and transitive on $M\setminus\{0\}$, and either $G^{(\infty)}\leqslant K$ or $K\leqslant{\mathbf{Z}}(G^{(\infty)})$. Suppose that $G^{(\infty)}\leqslant K$. Then $G/K$ is isomorphic to a quotient group of $G/G^{(\infty)}\lesssim{\Gamma\mathrm{L}}(1,2^{f})$. So
$|G/K|\leqslant|{\Gamma\mathrm{L}}(1,2^{f})|=f(2^{f}-1)\leqslant\frac{n}{2}(2^{\frac{n}{2}}-1)<2^{n}-1=|M|-1,$
which is a contradiction since $G/K$ is intransitive on $M\setminus\{0\}$.
We thus have that $K\leqslant{\mathbf{Z}}(G^{\infty})$. So $G/K$ is faithful and transitive on $M\setminus\{0\}$, and has a simple composition factor isomorphic to one of ${\mathrm{PSL}}_{m}(q)$ ($m\geqslant 2$ and $(m,f)\neq(2,1)$), ${\mathrm{PSp}}_{2m}(q)$ ($m\geqslant 2$ and $(m,q)\neq(2,2)$) and ${\mathrm{G}}_{2}(q)$ ($q\geqslant 4$) for even $q$. Note that none of these groups is isomorphic to ${\mathrm{A}}_{6}$, ${\mathrm{A}}_{7}$ or ${\mathrm{PSU}}_{3}(3)$, inspecting the candidates for $G/K$ in Theorem 2.4, we conclude that $(G/K)^{(\infty)}\cong G^{(\infty)}$. It follows that $K\cap G^{(\infty)}=1$ and $G^{(\infty)}$ is transitive on $M\setminus\{0\}$, and further $G^{(\infty)}$ is also transitive on $V\setminus\{0\}$.
Let $U$ be the natural ${\mathbb{F}}_{q}G^{(\infty)}$-module, and let $\phi:x\mapsto x^{2}$ be the Frobenius automorphism of ${\mathbb{F}}_{q}$. Clear that $U_{{\mathbb{F}}_{2}}\cong V$. By Lemma 2.2, $\Lambda^{2}_{{\mathbb{F}}_{2}}(V)$ is a direct sum of ${\mathbb{F}}_{2}G^{(\infty)}$-submodules $A$ and $B_{j}$’s for $1\leqslant j\leqslant\lfloor\frac{f}{2}\rfloor$, where
  1. (i)
    $A\cong\Lambda_{{\mathbb{F}}_{q}}^{2}(U)_{{\mathbb{F}}_{2}}$;
  2. (ii)
    $B_{j}\cong(U\otimes U^{\phi^{j}})_{{\mathbb{F}}_{2}}$ for $1\leqslant j<\frac{f}{2}$;
  3. (iii)
    when $f$ is even, $2\cdot B_{f/2}\cong(U\otimes U^{\phi^{f/2}})_{{\mathbb{F}}_{2}}$.
Lemma 2.6 deduces that each $B_{j}$ is an irreducible ${\mathbb{F}}_{2}G^{(\infty)}$-module. Suppose that $B_{j}$ is not a subspace of $W$ for some $j$. Then $M\cong B_{j}$ as ${\mathbb{F}}_{2}G^{(\infty)}$-modules. It follows that $\dim M=\dim B_{j}$. Note that $\dim B_{j}=n^{2}/f$ for $j\neq f/2$ and $\dim B_{f/2}=n^{2}/(2f)$ when $f$ is even. Hence, $M\cong B_{j}$ only when $j=f/2$, $f=n/2$ and $G^{(\infty)}\cong{\mathrm{SL}}_{2}(2^{f})$. In this case, the induced linear group of $G^{(\infty)}$ on $B_{f/2}$ is isomorphic to $\Omega^{-}_{4}(2^{f/2})$ (see [14, page 45]), which cannot be transitive on non-zero vectors of ${\mathbb{F}}_{2}^{2f}$. Thus, $B_{j}\leqslant W$ for each $j$, and then $M$ is ${\mathbb{F}}_{2}G^{(\infty)}$ isomorphic to a quotient of $A\cong(\Lambda^{2}_{{\mathbb{F}}_{q}}(U))_{{\mathbb{F}}_{2}}$. Suppose that $G^{(\infty)}\cong{\mathrm{Sp}}_{2m}(q)$ for $m\geqslant 2$ (or $G^{(\infty)}\cong G_{2}(q)$) with $q=2^{f}$. Let $T$ and $T_{0}$ be ${\mathbb{F}}_{q}G^{(\infty)}$-submodules of $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ give in Lemma 2.8. Assume that $m$ is even. Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is indecomposable with $T$ a maximal ${\mathbb{F}}_{q}G^{(\infty)}$-submodule by Lemma 2.8. Note that $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)/T$ is a trivial ${\mathbb{F}}_{q}G^{(\infty)}$-module. This yields that $M$ is a trivial ${\mathbb{F}}_{2}G^{(\infty)}$-module, a contradiction. Assume that $m$ is odd. Then $T$ is irreducible and cannot be written over any subfield by Lemma 2.8. Hence, $T_{{\mathbb{F}}_{2}}$ is an irreducible ${\mathbb{F}}_{2}G^{(\infty)}$-module. This deduces that $M\cong S_{{\mathbb{F}}_{2}}$ as ${\mathbb{F}}_{2}T$-modules. Note that $\dim T_{{\mathbb{F}}_{2}}=f\cdot(\binom{n/f}{2}-1)$. By some elementary calculations, we obtain that $\dim T_{{\mathbb{F}}_{2}}=\dim M=n$ has no no integer solutions.
Suppose that $G^{(\infty)}\cong{\mathrm{SL}}_{m}(q)$ for some $m\geqslant 2$ and $q=2^{f}$. Lemma 2.7 says $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is an irreducible ${\mathbb{F}}_{q}G^{(\infty)}$-module and cannot be written over any subfield. Hence, $(\Lambda^{2}_{{\mathbb{F}}_{q}}(U))_{{\mathbb{F}}_{2}}$ is an irreducible ${\mathbb{F}}_{q}G_{0}$-module and $M\cong(\Lambda^{2}_{{\mathbb{F}}_{q}}(U))_{{\mathbb{F}}_{2}}$. Note that $\dim(\Lambda^{2}_{{\mathbb{F}}_{q}}(U))_{{\mathbb{F}}_{2}}=f\cdot\binom{n/f}{2}$ is equal to $\dim M=n$. We obtain that $f=n/3$, $m=3$ and $G^{(\infty)}\cong{\mathrm{SL}}_{3}(2^{n/3})$. In this case, $M\cong A\cong\Lambda^{2}_{{\mathbb{F}}_{q}}(U)_{{\mathbb{F}}_{2}}$ is isomorphic to the dual of $V$ deduced by Lemma 2.7. ∎
3. Proof of Theorem 1.1
For a non-abelian special $p$-group $N$, the automorphism group $A={\mathrm{Aut}}(N)$ naturally acts on the characteristic subgroup ${\mathbf{Z}}(N)$ and induces an action on the factor group $N/{\mathbf{Z}}(N)$. Viewing the elementary abelian $p$-groups $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$ as vector spaces over the field ${\mathbb{F}}_{p}$, the actions of $A$ on them can be viewed as linear actions, and these give rise to two natural ${\mathbb{F}}_{p}A$-modules, denoted by $V$ and $M$, respectively. The following lemma enables us to employ methods from group representation theory to study special $p$-groups.
Lemma 3.1.
Let $N$ be a non-abelian special $p$-group with $p$ prime such that $N/{\mathbf{Z}}(N)\cong{\mathbb{Z}}_{p}^{n}$. Let $K$ and $L$ be the kernels of $A:={\mathrm{Aut}}(N)$ acting on $V:=N/{\mathbf{Z}}(N)$ and $M:={\mathbf{Z}}(N)$, respectively. Let $G=A/K$. Then the following statements hold:
  1. (i)
    [9, Lemma 3] $K$ is an elementary abelian $p$-group of order $|{\mathbf{Z}}(N)|^{n}$;
  2. (ii)
    $K\leqslant L$, and then both $V$ and $M$ are natural ${\mathbb{F}}_{p}G$-modules;
  3. (iii)
    the number of $A$-orbits on $N$ is equal to $o(V)+o(M)-1$ where $o(V)$ and $o(M)$ are numbers of $G$-orbits on $V$ and $M$ respectively;
  4. (iv)
    $M\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ for some ${\mathbb{F}}_{p}G$-submodule $W$ of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$.
Proof.
Let $C={\mathbf{Z}}(N)=\Phi(N)=N^{\prime}$ for convenience. By definition, $V:=N/C$ is a faithful ${\mathbb{F}}_{p}G$-module as $K$ is the kernel of $G$ on $V$.
Suppose that $N=\langle x_{1},...,x_{n}\rangle$. Then $M=N^{\prime}$ is generated by commutators $[x_{i},x_{j}]$ for $i,j=1,...,n$. By definition, each element $\sigma\in K$ fixes all of the cosets ${\mathbf{Z}}(N)x_{i}$, with $1\leqslant i\leqslant n$, so
$x_{i}^{\sigma}=a_{i}x_{i}$, where $a_{i}\in{\mathbf{Z}}(N)$ for each $i=1,...,n$.
Hence, $[x_{i},x_{j}]^{\sigma}=[a_{i}x_{i},a_{j}x_{j}]=[x_{i},x_{j}]$ as $a_{i},a_{j}\in{\mathbf{Z}}(N)$. Thus $K$ acts trivially on $M$, and so $K\leqslant L$. It follows that $G=A/K$ acts on $M$ with kernel $L/K$, which yields that $M$ is an ${\mathbb{F}}_{p}G$-module, as in (ii).
Note that $A$ normalizes $C$, and partitions $N$ into two parts $C$ and $N\setminus C$. Thus, the number of $A$-orbits on $N$ is equal to the sum of numbers of $A$-orbits on $C$ and on $N\setminus C$. As usual, let $\sim$’ denote conjugation relation.
First, as $K\leqslant L$ and $L$ fixes every element of $C$, for any elements $a,b\in C=M$,
$a\sim b$ under $A$ $\Longleftrightarrow a\sim b$ under $A/L$ $\Longleftrightarrow a\sim b$ under $A/K=G$.
Thus the $A$-orbits on $C$ equal the $G$-orbits on $M=C$, and hence the number of $A$-orbits on $C$ is equal to $o(M)$.
Next, as $K$ fixes each coset $Cx_{i}$, for each $\alpha\in K$, there exist $c_{1},\dots,c_{n}\in C$ such that
$(x_{1},\dots,x_{n})^{\alpha}=(c_{1}x_{1},\dots,c_{n}x_{n}).$
Thus $\alpha$ is uniquely determined by the tuple $(c_{1},\dots,c_{n})$, and so
$|K|\leqslant|\{(c_{1},\dots,c_{n})\mid c_{i}\in C\}|=|C|^{n}.$
By Lemma 3.1 (i), we have $|K|=|C|^{n}$, and thus $|K|=|\{(c_{1},\dots,c_{n})\mid c_{i}\in C\}|$. Hence elements of $K$ and the tuples $(c_{1},\dots,c_{n})$ with $c_{i}\in C$ are one-to-one correspondent. It follows that elements of the same coset $Cx$ are conjugate under $K$. Hence, two elements $x,y\in N\setminus C$ are conjugate under ${\mathrm{Aut}}(N)$ if and only if the cosets $Cx,Cy$ are conjugate under $G=A/K$. Equivalently, the number of $A$-orbits on $N\setminus C$ is equal to the number of $G$-orbits on non-identity elements of $N/C=V$, which equals $o(V)-1$. Thus, the number of $A$-orbits on $N$ is equal to $o(V)+o(M)-1$, as in part (iii). For $x,y,z\in N$, we notice that $[x,y]=[y,x]^{-1}$ and $[x,yz]=[x,y]^{z}[x,z]=[x,y][x,z]$ as $N^{\prime}={\mathbf{Z}}(N)$. Hence the commutator operator induces an alternative linear map:
$\varphi:\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=V\wedge V\rightarrow M$ such that $\varphi:xC\wedge yC\mapsto[x,y]$.
For any element $g\in G$, we have that
$(xC\wedge yC)^{g\varphi}=[(xC)^{g},(yC)^{g}]=[x,y]^{g}=(xC\wedge yC)^{\varphi g}.$
Thus $\varphi$ is an ${\mathbb{F}}_{p}G$-homomorphism from $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ to $M$, and $M$ is isomorphic to $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ for some ${\mathbb{F}}_{p}G$-submodule $W$ of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$, as in (iv). ∎
Now we are ready to produce a list of finite $p$-groups with a single fusion class of elements of order $p$, and prove Theorem 1.1.
Proof of Theorem 1.1.
Let $N$ be a $p$-group of which subgroups of order $p$ are all fused.
If $N$ is abelian, then obviously $N$ is homocyclic. If $p$ is an odd prime, then $N$ is a homocyclic group by Shult [22]. If $p=2$ and $N$ has only one involution, then $N$ is a generalized quaternion group.
We thus assume that $N$ is a non-abelian $2$-group which has more than one involution, and ${\mathrm{Aut}}(N)$ is transitive on the involutions. By Gross [10], one of the following holds:
  1. (i)
    $N$ has exponent $4$ and nilpotence class $2$, $|N|=|{\mathbf{Z}}(N)|^{2}$ or $|{\mathbf{Z}}(N)|^{3}$, and $N^{\prime}=N^{2}={\mathbf{Z}}(N)$;
  2. (ii)
    $N$ has exponent $8$ with nilpotence class $3$.
By Bryukhanova [3], a group $N$ in case (i) with $|N|=|{\mathbf{Z}}(N)|^{3}$ indeed has ${\mathrm{Aut}}(N)$ solvable, and is a Suzuki $2$-group. Wilkens [24] excluded groups in case (ii), that is, the nilpotence class of $N$ is not equal to $3$.
To complete the proof, we may assume that $N$ is a non-abelian special $2$-group of order equal to $|{\mathbf{Z}}(N)|^{2}$.
Then $\Phi(N)=N^{\prime}N^{2}=N^{\prime}={\mathbf{Z}}(N)$, and $N$ is a special $2$-group. Let $V$ and $M$ be the vector space corresponding to $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$ respectively. Let $G={\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$, the group induced by ${\mathrm{Aut}}(N)$ on the factor group $N/\Phi(N)$. Lemma 3.1 shows that $V$ and $M$ are ${\mathbb{F}}_{2}G$-modules for $G$, and $M$ is ${\mathbb{F}}_{2}G$-isomorphic to a quotient of $\Lambda_{{\mathbb{F}}_{2}}^{2}(V)$. First, we show that $G$ acts transitively on non-zero vectors of both $V$ and $M$. Recall that ${\mathbf{Z}}(N)=N^{\prime}$ consists of all involutions of $N$ and the identity. Hence, non-zero vectors of $M$ are in the same orbits of ${\mathrm{Aut}}(N)$. Since $G^{{\mathbf{Z}}(N)}={\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$ by Lemma 3.1(ii), it follows that $G$ is transitive on non-zero vectors of $M$. Note that every element in $N\setminus{\mathbf{Z}}(N)$ is of order $4$. Suppose that $x,y\in N\setminus{\mathbf{Z}}(N)$. Then $x^{2}$ and $y^{2}$ are involutions, and hence there exists $\alpha\in{\mathrm{Aut}}(N)$ such that $(x^{2})^{\alpha}=y^{2}$. This yields that $(x^{\alpha})^{2}=y^{2}$. We may decompose $N=x_{1}{\mathbf{Z}}(N)\cup x_{2}{\mathbf{Z}}(N)\cup\cdots\cup x_{|{\mathbf{Z}}(N)|}{\mathbf{Z}}(N)$ for some $x_{i}\in N\setminus{\mathbf{Z}}(N)$ as $|N/{\mathbf{Z}}(N)|=|{\mathbf{Z}}(N)|$. Then, for each representative $x_{i}$ and any element $a\in M$, we have $(ax_{i})^{2}=a^{2}x_{i}^{2}=x_{i}^{2}$. Thus, we have
[…]
$\Longleftrightarrow x^{\alpha}{\mathbf{Z}}(N)=y{\mathbf{Z}}(N)\Longleftrightarrow(x{\mathbf{Z}}(N))^{\alpha}=y{\mathbf{Z}}(N).$
Hence, any two non-identity elements of $V:=N/{\mathbf{Z}}(N)$ are in the same orbits of ${\mathrm{Aut}}(N)$. Thus, $G$ is transitive on non-zero vectors of $V$.
Suppose that ${\mathrm{Aut}}(N)$ is non-solvable. Then $G$ is non-solvable by Lemma 3.1(i). Since $\dim M=\dim V$, we have that $G^{(\infty)}\cong{\mathrm{SL}}_{3}(2^{f})$ and $M$ is ${\mathbb{F}}_{2}G^{(\infty)}$-isomorphic to the dual $V$ by Theorem 1.2. Let $v=x{\mathbf{Z}}(N)$ be a non-zero vector of $V$ where $x\in N\setminus{\mathbf{Z}}(N)$. Then $x$ is of order $4$, and $w:=x^{2}$ is a non-zero vector in $M$. Note that an element of $G^{(\infty)}$ stabilizing $v$ must stabilizes $w$ since $(ax)^{2}=x^{2}$ for any $a\in{\mathbf{Z}}(N)$, and hence $G^{(\infty)}_{v}\leqslant G^{(\infty)}_{w}$, which contradicts the fact that $M$ is the dual ${\mathbb{F}}_{2}G^{(\infty)}$-module of $V$ for $G^{(\infty)}\cong{\mathrm{SL}}_{3}(2^{f})$. Thus, ${\mathrm{Aut}}(N)$ is solvable, and $N$ is a Suzuki $2$-group by Shaw’s Theorem [21]. The solvability of ${\mathrm{Aut}}(N)$ is proved by Bryukhanova in [3, Theorem 1.2]. ∎
4. Finite $2$-groups with three fusion classes
In this section, we introduce the three families of AT-$2$-groups, and then prove Corollary 1.3. The first family is $A_{2}(n,\theta)$, a Suzuki $2$-group of type A, see [11, page 82].
Definition 4.1.
Let $\theta\in{\mathrm{Gal}}({\mathbb{F}}_{2^{n}}/{\mathbb{F}}_{2})$ such that $|\theta|\neq 1$ is odd. For $a,b\in{\mathbb{F}}_{2^{n}}$, define
$A_{2}(n,\theta)=\langle M(a,b,\theta)\mid a,b\in{\mathbb{F}}_{2^{n}}\rangle,$
where $M(a,b,\theta)$ is a matrix of the form
$M(a,b,\theta)=\begin{pmatrix}1&a&b\\ 0&1&a^{\theta}\\ 0&0&1\end{pmatrix}.$
We note that, if $n=2k+1$ and $x^{\theta}=x^{k+1}$, then $A_{2}(2k+1,\theta)$ is isomorphic to a Sylow $2$-subgroup of the Suzuki simple group ${\mathrm{Sz}}(2^{2k+1})$, see [6, Proposition 13.4(vi)]. For any odd integers $n\neq 1$, the groups $A_{2}(n,\theta)$ are AT-groups by Zhang in [25, Lemma 2.2].
The group $N=A_{2}(n,\theta)$ is a special $2$-group with ${\mathbf{Z}}(N)=\langle M(0,b,\theta)\mid b\in{\mathbb{F}}_{2^{n}}\rangle\cong{\mathbb{Z}}_{2}^{n}$. The matrix $D=\mathrm{diag}(\lambda^{-1},1,\lambda^{\theta})$ induces an automorphism $\xi$ of $N$, that is
$\xi:M(a,b,\theta)\mapsto D^{-1}M(a,b,\theta)D=M(a\lambda,b\lambda\lambda^{\theta},\theta),$
and $|\xi|=2^{n}-1$. The Frobenius automorphism $\phi$ of ${\mathbb{F}}_{2}^{2^{n}}$ induces an automorphism of $A_{2}(n,\theta)$ of order $n$, and $\langle\xi,\phi\rangle={\Gamma\mathrm{L}}_{1}(2^{n})$.
Lemma 4.2.
The group $N=A_{2}(n,\theta)$ is an AT-group, and ${\mathrm{Aut}}(N)\cong 2^{n^{2}}{:}{\Gamma\mathrm{L}}_{1}(2^{n})$.
Proof.
Obviously, $M(a,b,\theta)$ and $M(a\lambda,b\lambda\lambda^{\theta},\theta)$ do not belong to the same coset of ${\mathbf{Z}}(N)$. It follows that $\langle\xi\rangle$ acts on $N/{\mathbf{Z}}(N)$ faithfully, and hence $\langle\xi,\phi\rangle\cong{\Gamma\mathrm{L}}_{1}(2^{n})$ acts on $N/{\mathbf{Z}}(N)$ faithfully. Let $K$ be the kernel of ${\mathrm{Aut}}(N)$ acting on $N/{\mathbf{Z}}(N)$. Lemma 3.1(i) deduces that $K\cong{\mathbb{Z}}_{2}^{n^{2}}$, and thus
$\langle K,\xi,\phi\rangle\cong 2^{n^{2}}{:}{\Gamma\mathrm{L}}_{1}(2^{n})\leqslant{\mathrm{Aut}}(N)$.
Noting that ${\mathrm{Aut}}(N)$ is solvable by Theorem 1.1(iii), we have that ${\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$ is isomorphic to a solvable subgroup of ${\mathrm{GL}}_{n}(2)$ which is transitive on the non-zero vectors of ${\mathbb{F}}_{2}^{n}$. By Theorem 2.4, we obtain that $N/K={\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}\lesssim{\Gamma\mathrm{L}}_{1}(2^{n})$, and so ${\mathrm{Aut}}(N)\cong 2^{n^{2}}{:}{\Gamma\mathrm{L}}_{1}(2^{n})$. ∎
Sylow $2$-subgroups of ${\mathrm{SU}}_{3}(2^{n})$ are Suzuki $2$-group of type B (see [12, page 314]). Equip the space ${\mathbb{F}}_{q^{2}}^{3}$ ($q=2^{n}$) with the unitary form:
$\bigl{(}(x_{1},x_{2},x_{3}),(y_{1},y_{2},y_{3})\bigr{)}=x_{1}y_{3}^{q}+x_{2}y_{2}^{q}+x_{3}y_{1}^{q}.$
Then the following group $B_{2}(n)$ is a Sylow $2$-subgroup of ${\mathrm{SU}}_{3}(q)$, see [7, page 249].
Definition 4.3.
Let $q=2^{n}$, and $B_{2}(n)=\langle M(a,b)\mid a,b\in{\mathbb{F}}_{q^{2}},b+b^{q}+aa^{q}=0\rangle$, where
$M(a,b)=\begin{pmatrix}1&a&b\\ 0&1&a^{q}\\ 0&0&1\end{pmatrix}.$
It was noticed in Zhang [25, Lemma 2.3] that $B_{2}(n)$ is an AT-group, with ${\mathbf{Z}}(N)=\langle M(0,b)\mid b\in{\mathbb{F}}_{q^{2}},b+b^{q}=0\rangle\cong{\mathbb{Z}}_{2}^{n}$.
Lemma 4.4.
Let $N=B_{2}(n)$. Then $N$ is an AT-group, and ${\mathrm{Aut}}(N)\cong 2^{2n^{2}}{:}{\Gamma\mathrm{L}}_{1}(2^{2n})$.
Proof.
Let $D=\mathrm{diag}(\lambda^{-q},\lambda^{1-q},\lambda)$, where $\langle\lambda\rangle={\mathbb{F}}_{q^{2}}^{\times}$. Then $D^{-1}M(a,b)D=M(q\lambda,b\lambda^{q+1})$, and $b\lambda^{q+1}+(b\lambda^{q+1})^{q}+(a\lambda)(a\lambda)^{q}=\lambda^{1+q}(b+b^{q}+a^{1+q})=0$. Thus, $D$ induces an automorphism of $N$, denoted by $\xi$. Let $\phi$ be the Frobenius automorphism $x\mapsto x^{2}$ on ${\mathbb{F}}_{q^{2}}$, which induces an automorphism of $N$. As $M(a,b)$ and $M(q\lambda,b\lambda^{q+1})$ are not in the same coset of ${\mathbf{Z}}(N)$, it follows that $\langle\xi,\phi\rangle\cong{\Gamma\mathrm{L}}_{1}(2^{2n})$ acts faithfully on $N/{\mathbf{Z}}(N)$.
Arguing as in the proof of Lemma 4.2, the kernel $K$ of ${\mathrm{Aut}}(N)$ on $N/{\mathbf{Z}}(N)$ is isomorphic to ${\mathbb{Z}}_{2}^{n^{2}}$, and ${\mathrm{Aut}}(N)/K=\langle\xi,\phi\rangle$. Therefore, we conclude that ${\mathrm{Aut}}(N)\cong 2^{2n^{2}}{:}{\Gamma\mathrm{L}}_{1}(2^{2n})$, and $N$ is an AT-group. ∎
The following Suzuki $2$-group was first given in case (viii) of the main Theorem in [8].
Definition 4.5.
For a generator $\epsilon$ of ${\mathbb{F}}_{2^{6}}^{\times}$, define
$P(\epsilon)=\{(a,x)\in{\mathbb{F}}_{2^{6}}\times{\mathbb{F}}_{2^{3}}\mid(a,x)(b,y)=(a+b,x+y+ab^{2}\epsilon+a^{8}b^{16}\epsilon^{8})\}.$
The next lemma shows that the definition of $P(\epsilon)$ is independent of the choice of $\epsilon$.
Lemma 4.6.
For any two generators $\epsilon$ and $\lambda$ of ${\mathbb{F}}_{2^{6}}^{\times}$, we have $P(\epsilon)\cong P(\lambda)$.
Proof.
Let $\epsilon$ be a generator of ${\mathbb{F}}_{2^{6}}^{\times}$. Let ${\mathrm{Tr}}:x\mapsto x+x^{8}$ be the trace map on ${\mathbb{F}}_{2^{6}}$ over ${\mathbb{F}}_{2^{3}}$. For convenience, let $f_{\epsilon}(a)={\mathrm{Tr}}(a^{3}\epsilon)$ and let $g_{\epsilon}(a,b)={\mathrm{Tr}}\bigl{(}(a+b)ab\epsilon\bigr{)}$. Calculation shows
$(a,x)^{2}=(0,f_{\epsilon}(a)),\mbox{ and }[(a,x),(b,y)]=\bigl{(}0,g_{\epsilon}(a,b)\bigr{)}\mbox{, in $P(\epsilon)$}.$
Let $X=(x_{1},...,x_{6})$ be a basis of ${\mathbb{F}}_{2^{6}}^{+}$, and $Z=(z_{1},z_{2},z_{3})$ be a basis of ${\mathbb{F}}_{2^{3}}^{+}$. Then the generating relations of the group $P(\epsilon)$ can be obtained by the coefficients of $f_{\epsilon}(x_{i})$ and $g_{\epsilon}(x_{i},x_{j})$ with respect to the basis $Z$ of ${\mathbb{F}}_{2^{3}}^{+}$ for each $i,j=1,...,6$.
First, we show that $P(\epsilon)\cong P(\epsilon^{3k+1})$ for any generator $\epsilon$ and any integer $k$. Fix $Z=(1,\epsilon^{9},\epsilon^{18})$. Set $x_{i}=\epsilon^{i+k-1}$ and $y_{i}=\epsilon^{i-1}$ for $i=1,...,6$. Then both $X=(x_{1},...,x_{6})$ and $Y=(y_{1},...,y_{6})$ are basis of ${\mathbb{F}}_{2^{6}}^{+}$. Note that
$f_{\epsilon^{3k+1}}(a)$ $={\mathrm{Tr}}(a^{3}\epsilon^{3k+1})={\mathrm{Tr}}((a\epsilon^{k})^{3}\epsilon)=f_{\epsilon}(a\epsilon^{k}),$
$g_{\epsilon^{3k+1}}(a,b)$ $={\mathrm{Tr}}\bigl{(}(a+b)ab\epsilon^{3k+1}\bigr{)}={\mathrm{Tr}}\bigl{(}((a\epsilon^{k})+(b\epsilon^{k}))(a\epsilon^{k})(b\epsilon^{k})\bigr{)}=g_{\epsilon}(a\epsilon^{k},b\epsilon^{k}).$
Thus, under the basis $Z$, the coefficients of $f_{\epsilon}(x_{i})$ and $f_{\epsilon^{3k+1}}(y_{i})$ are equal, and the coefficients of $g_{\epsilon}(x_{i},x_{j})$ and $g_{\epsilon^{3k+1}}(y_{i},y_{j})$ are equal for $1\leqslant i,j\leqslant 6$. Hence $P(\epsilon)\cong P(\epsilon^{3k+1})$ for each integer $k$.
We thus conclude that, for a given generator $\epsilon$, if $s\equiv t\pmod{3}$ then $P(\epsilon^{s})\cong P(\epsilon^{t})$. Note that $\epsilon^{3k}$ is not an generator of ${\mathbb{F}}_{2^{6}}^{\times}$ for any integer $k$. Hence, for any generator $\lambda$ of ${\mathbb{F}}_{2^{6}}^{\times}$, $P(\lambda)$ is isomorphic to one of $P(\epsilon)$ and $P(\epsilon^{2})$.
Choose $X=(1,\epsilon,...,\epsilon^{5})$ and $Z=(1,\epsilon^{9},\epsilon^{18})$. It is not hard to see that the coefficients of $f_{\epsilon}(x_{i})$ and $g_{\epsilon}(x_{i},x_{j})$ with respect basis $Z$ only depend on the minimal polynomial of $\epsilon$ over ${\mathbb{F}}_{2}$. Notice that the minimal polynomials of $\epsilon$ and $\epsilon^{2}$ are the same since $x\mapsto x^{2}$ is the Frobenius automorphism of ${\mathbb{F}}_{2^{6}}$. Thus we have $P(\epsilon)\cong P(\epsilon^{2})$.
Therefore, $P(\lambda)$ is isomorphic to $P(\epsilon)$ for any generators $\lambda$ and $\epsilon$ of ${\mathbb{F}}_{2^{6}}^{\times}$. ∎
The automorphism group of $P(\epsilon)$ can be computed by Magma, led to the next lemma.
Lemma 4.7.
Let $N=P(\epsilon)$ for a generator $\epsilon$ of ${\mathbb{F}}_{2^{6}}^{\times}$. Then $N$ is an AT-group with ${\mathrm{Aut}}(N)\cong 2^{18}{:}(7{:}9)$.
We remark that some generators $\epsilon$ of ${\mathbb{F}}_{2^{6}}^{\times}$ satisfy minimal polynomial $\epsilon^{6}+\epsilon^{4}+\epsilon^{3}+\epsilon+1$, and with this polynomial we may write the generating relations for $P(\epsilon)$ explicitly, which makes computation with $P(\epsilon)$ in Magma efficiently:
$P(\epsilon)=$ $\langle x_{i},z_{j}\mbox{ for $1\leqslant i\leqslant 6$ and $1\leqslant j\leqslant 3\rangle$, where}$
$\begin{array}[]{ll}&z_{j}^{2}=[x_{i},z_{j}]=[z_{k},z_{\ell}]=1\mbox{ for each $i,j,k,\ell$,}\\ &x_{1}^{2}=x_{3}^{2}=z_{2},\ x_{2}^{2}=z_{2}z_{3},\ x_{4}^{2}=z_{3},\ x_{5}^{2}=z_{1}z_{2}z_{3},\ x_{6}^{2}=z_{3},\\ &[x_{1},x_{2}]=[x_{3},x_{5}]=[x_{3},x_{6}]=z_{1}z_{2},\ [x_{1},x_{3}]=z_{1}z_{3},\ [x_{1},x_{4}]=z_{3},\\ &[x_{1},x_{5}]=[x_{3},x_{4}]=[x_{5},x_{6}]=z_{2},\ [x_{1},x_{6}]=1,\ [x_{2},x_{6}]=z_{1}z_{2}z_{3},\\ &[x_{2},x_{3}]=[x_{2},x_{4}]=[x_{4},x_{6}]=z_{1},\ [x_{2},x_{5}]=[x_{4},x_{5}]=z_{2}z_{3}\rangle.\end{array}$
Dornhoff [8] determined finite groups $N$ of which a solvable group of automorphisms has exactly three orbits on $N$. In his classification, desired non-abelian $2$-groups are precisely $A_{2}(n,\theta)$, $B_{p}(n)$ and $P(\epsilon)$. Then the following lemma immediately holds since each Suzuki $2$-group has a solvable automorphism group by Theorem 1.1(iii).
Lemma 4.8.
If $N$ is a finite non-abelian $2$-group such that ${\mathrm{Aut}}(N)$ has exactly three orbits on $N$, then $N=A_{2}(n,\theta)$, $B_{p}(n)$ or $P(\epsilon)$.
Now we are ready to prove Corollary 1.3. Recall that two elements are conjugate under the automorphism group are said to be fused.
Proof of Corollary 1.3.
First, let $N=A_{2}(n,\theta)$, $B_{p}(n)$ or $P(\epsilon)$, as in part (iv) of Corollary 1.3. By Lemmas 4.2(ii), 4.4(ii) and 4.7, the group $N$ is an AT-group, as in part (iii) of Corollary 1.3. This yields (iv)$\Rightarrow$(iii)$\Rightarrow$(ii)$\Rightarrow$(i), as by definition an AT-group is an FIF-group, and cyclic subgroups of an FIF-group $N$ of the same order are fused.
Now assume that cyclic subgroups of $N$ of the same order are fused. Then involutions of $N$ are fused, and $N$ is a Suzuki $2$-group. Obviously, non-identity elements of ${\mathbf{Z}}(N)$ are all involutions, and elements in $N\setminus{\mathbf{Z}}(N)$ are all of order $4$. By definition, for any elements $x,y\in N\setminus{\mathbf{Z}}(N)$, there exists $\alpha\in{\mathrm{Aut}}(N)$ such that $x^{\alpha}=y$ or $x^{\alpha}=y^{-1}$. Hence, $(x{\mathbf{Z}}(N))^{\alpha}=y{\mathbf{Z}}(N)$, and ${\mathrm{Aut}}(N)$ is transitive on non-identity elements of $N/{\mathbf{Z}}(N)$. By Lemma 3.1(iii), ${\mathrm{Aut}}(N)$ has exactly three orbits on $N$, and thus
  1. (a)
    elements of the same order are in the same orbit of ${\mathrm{Aut}}(N)$, and $N$ is an AT-group, and also an FIF-group;
  2. (b)
    $N=A_{2}(n,\theta)$, $B_{p}(n)$ or $P(\epsilon)$ by Lemma 4.8.
That is to say, part (i) of Corollary 1.3 implies parts (ii)-(iv). This deduces that parts (i)-(iii) of Corollary 1.3 are equivalent as obviously (iii)$\Rightarrow$(ii)$\Rightarrow$(i). So all the four statements of Corollary 1.3 are equivalent, completing the proof. ∎

### v2

@@
 Building upon previous results, a classification is given of finite $p$-groups whose elements of order $p$ are all fused.
-This particularly confirms a conjecture of Gross proposed in 1976 on $2$-automorphic $2$-groups, which are $2$-groups with involutions forming a single fusion classes.
+This particularly confirms a conjecture of Gross proposed in 1976 on $2$-automorphic $2$-groups, which are $2$-groups with involutions forming a single fusion class.
 As a consequence, two open problems regarding AT-groups and FIF-groups are solved.
@@
 Suzuki $2$-groups; $2$-automorphic $2$-groups; exterior squre; dual module
+2020 Mathematics Subject Classification:
+20D15, 20D45, 20C33.
 1. Introduction
-It is well-known that finite $p$-groups with a single subgroup of order $p$ are cyclic or a generalized quaternion group.
+It is well-known that finite $p$-groups with a single subgroup of order $p$ are cyclic or a generalized quaternion group [11, page 199].
 The class of finite $p$-groups with a single fusion class of elements of order $p$ has been extensively studied.
 (Recall that a fusion class of a group $N$ is an orbit of ${\mathrm{Aut}}(N)$ on the elements.) The study was initiated by G.
-Higman 1963 when he classified finite non-abelian $2$-groups (with more than one involution) that has a cyclic group of automorphisms transitive on the involutions, called Suzuki $2$-groups.
+Higman [13] in 1963 when he classified finite non-abelian $2$-groups (with more than one involution) that have cyclic groups of automorphisms transitive on the involutions, called Suzuki $2$-groups.
 For odd primes $p$, Shult [24] showed that homocyclic $p$-groups are the only $p$-groups of which elements of order $p$ form a single fusion class.
 Shaw [23] showed that a finite non-abelian $2$-group (with more than one involution) which has a solvable group of automorphisms transitive on the involutions is a Suzuki $2$-group.
-Higman’s classification associated with Shaw’s work forms Sections 5-7 of Huppert’s book [12, Chapter VIII].
+Higman’s classification associated with Shaw’s work forms Sections 5-7 of Chapter VII of the book written by Huppert and Blackburn [14].
 In 1976, Gross [12] studied finite $2$-groups with a single fusion class of involutions, called $2$-automorphic $2$-groups.
@@
     $N$ is a Suzuki $2$-group, and ${\mathrm{Aut}}(N)$ is solvable with a cyclic subgroup acting transitively on the set of involutions.
+We use $G^{(\infty)}$ for the unique perfect group appearing in the derived series of $G$.
 The proof of Theorem 1.1 for $2$-groups depends on certain representations of finite transitive linear groups over ${\mathbb{F}}_{2}$, stated below, which is independently interesting.
@@
 A group is called an FIF-group if any two elements of the same order are fused or inverse-fused.
-Praeger and the first-named author [16] give a description of finite FIF-groups, with the problem of classifying FIF-$2$-groups left open, see [16, Problem 1.5].
+Praeger and the first-named author [18] give a description of finite FIF-groups, with the problem of classifying FIF-$2$-groups left open, see [18, Problem 1.5].
 As a corollary of Theorem 1.1, these problems are solved.
@@
     $N=A_{2}(n,\theta)$, $B_{2}(n)$, or $P(\epsilon)$; see Definitions 4.1, 4.3 and 4.5.
-We remark that the automorphism groups of the groups $A_{2}(n,\theta)$, $B_{2}(n)$ and $P(\epsilon)$ are determined in Lemmas 4.2, 4.4 and 4.7.
+The three families of groups in Corollary 1.3(iv) are exactly all $2$-groups with $3$ fusion classes, which were also obtained by Bors and Glasby in [2].
+We remark that the automorphism groups of the groups $A_{2}(n,\theta)$, $B_{2}(n)$ and $P(\epsilon)$ are determined in Lemmas 4.2, 4.4 and 4.7, respectively.
+The layout of the paper is as follows.
+In Section 2, we introduce some methods in representation theory and prove Theorem 1.2.
+We prove Theorem 1.1 in Section 3 by linking special $p$-groups with representation theory and using Theorem 1.2.
+The final section contains the proof of Corollary 1.3 together with the constructions of groups in the Corollary.
 Acknowledgments
-The authors acknowledge the support of NNSFC grant no. 11931005.
+The authors acknowledge the support of NNSFC grant no. 11931005. Moreover, they would like to thank the anonymous referee for their careful reading of the paper and valuable suggestions.
 2. Transitive linear groups and their exterior squares
@@
 If ${\mathrm{E}}$ is an extension field of ${\mathrm{F}}$, then $V^{\mathrm{E}}:=V\otimes{\mathrm{E}}$ is naturally an ${\mathrm{E}}G$-module.
-Some elementary properties of the operator $\cdot\otimes{\mathrm{E}}$ can be found in [12, Section VII.1], [1, Chapter 9] and [13, Chapter 9].
+Some elementary properties of the functor $-\otimes{\mathrm{E}}$ can be found in [14, Section VII.1], [1, Chapter 9] and [15, Chapter 9].
 For an ${\mathrm{F}}G$-module $V$ with representation $\rho$, we use the following two notations:
@@
   1. (i)
-    [12, Chapter VII, 1.16(a)] $(V_{\mathrm{K}})^{\mathrm{F}}\cong\sum_{\sigma\in{\mathrm{Gal}}({\mathrm{F}}/{\mathrm{K}})}V^{\sigma}$;
+    [14, Chapter VII, 1.16(a)] $(V_{\mathrm{K}})^{\mathrm{F}}\cong\sum_{\sigma\in{\mathrm{Gal}}({\mathrm{F}}/{\mathrm{K}})}V^{\sigma}$;
   2. (ii)
@@
   3. (iii)
-    [Deuring-Noether, see [12, Chapter VII, 1.22]] $V^{\mathrm{E}}\cong W^{\mathrm{E}}$ if and only if $V\cong W$.
+    [Deuring-Noether, see [14, Chapter VII, 1.22]] $V^{\mathrm{E}}\cong W^{\mathrm{E}}$ if and only if $V\cong W$.
 The exterior square $\Lambda^{2}_{{\mathrm{F}}}(V)$ of an ${\mathrm{F}}G$-module $V$ is $(V\otimes V)/S$ with $S=\langle v\otimes v:v\in V\rangle$.
@@
 Let $V={\mathbb{F}}_{2}^{n}$ be a vector space, and let $G\leqslant{\mathrm{GL}}(V)\cong{\mathrm{GL}}_{n}(2)$ be transitive on $V\setminus\{0\}$.
-These groups $G$ are known by the classification of the finite $2$-transitive permutation groups, and we list them in the next theorem, see [4, Table 7.3].
+These groups $G$ are known by the classification of the finite $2$-transitive permutation groups, and we list them in the next theorem, see [5, Table 7.3].
 Theorem 2.4.
@@
       ${\mathrm{G}}_{2}(2^{f})$ with $f\geqslant 2$.
-The following lemma can be directly deduced by Steinberg’s Twisted Tensor Product Theorem [23] (also see [14, Theorem 5.4.1]).
+The following lemma can be directly deduced by Steinberg’s Twisted Tensor Product Theorem [25] (also see [16, Theorem 5.4.1]).
 Lemma 2.5.
@@
 We record some classical results for exterior squares of natural modules for ${\mathrm{SL}}_{m}(q)$, ${\mathrm{Sp}}_{2m}(q)$ and ${\mathrm{G}}_{2}(q)$.
-Lemma 2.7 ([17, Theorem 2.2]).
+Lemma 2.7 ([19, Theorem 2.2]).
 Let $U={\mathbb{F}}_{q}^{m}$, and let $G={\mathrm{SL}}_{m}(q)$.
-Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is an irreducible ${\mathbb{F}}_{q}G$-module which cannot be written over an subfield of ${\mathbb{F}}_{q}$.
+Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is an irreducible ${\mathbb{F}}_{q}G$-module which cannot be written over any subfield of ${\mathbb{F}}_{q}$.
 In particular, if $m=3$, then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is isomorphic to the dual ${\mathbb{F}}_{q}G$-module of $V$.
-Results of exterior square of the natural modules of symplectic groups can be found in [5, 15, 19], and the result of the exceptional group of Lie type ${\mathrm{G}}_{2}(q)<{\mathrm{Sp}}_{6}(q)$ (even $q$) can be deduced from [20, Theorem 8.14].
+Results of exterior square of the natural modules of symplectic groups can be found in [6, 17, 21], and the result of the exceptional group of Lie type ${\mathrm{G}}_{2}(q)<{\mathrm{Sp}}_{6}(q)$ (even $q$) can be deduced from [22, Theorem 8.14].
 Lemma 2.8.
@@
 Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ has a maximal ${\mathbb{F}}_{q}G$-submodule $T$ of codimension $1$.
-Moerover, $T$ is indecomposable and one the following statements holds:
+Moreover, $T$ is indecomposable and one the following statements holds:
   1. (i)
@@
 Then $G$ is listed in Theorem 2.4.
-If $G$ is one of the small groups listed in Theorem 2.4(ii), then computations in Magma [2] shows that $\dim M=1$.
+If $G$ is one of the small groups listed in Theorem 2.4(ii), then computations in Magma [3] shows that $\dim M=1$, which is a contradiction as $\dim V=\dim M=1$ implies $G={\mathrm{GL}}_{1}(2)=1$.
 Thus, we only need to focus on the cases that $G^{(\infty)}\cong{\mathrm{SL}}_{m}(q)$ ($m\geqslant 2$ and $(m,f)\neq(2,1)$), ${\mathrm{Sp}}_{2m}(q)$ ($m\geqslant 2$ and $(m,q)\neq(2,2)$) or ${\mathrm{G}}_{2}(q)$ ($q\geqslant 4$) for $q=2^{f}$.
 Let $K$ be the kernel of $G$ acting on $M$.
-Then $G/K$ is faithful and transitive on $M\setminus\{0\}$, and either $G^{(\infty)}\leqslant K$ or $K\leqslant{\mathbf{Z}}(G^{(\infty)})$.
+Then $G/K$ is faithful and transitive on $M\setminus\{0\}$, and either $G^{(\infty)}\leqslant K$ or $K\cap G^{(\infty)}\leqslant{\mathbf{Z}}(G^{(\infty)})$.
 Suppose that $G^{(\infty)}\leqslant K$.
@@
 which is a contradiction since $G/K$ is intransitive on $M\setminus\{0\}$.
-We thus have that $K\leqslant{\mathbf{Z}}(G^{\infty})$.
+We thus have that $K\cap G^{(\infty)}\leqslant{\mathbf{Z}}(G^{(\infty)})$.
 So $G/K$ is faithful and transitive on $M\setminus\{0\}$, and has a simple composition factor isomorphic to one of ${\mathrm{PSL}}_{m}(q)$ ($m\geqslant 2$ and $(m,f)\neq(2,1)$), ${\mathrm{PSp}}_{2m}(q)$ ($m\geqslant 2$ and $(m,q)\neq(2,2)$) and ${\mathrm{G}}_{2}(q)$ ($q\geqslant 4$) for even $q$.
@@
 Hence, $M\cong B_{j}$ only when $j=f/2$, $f=n/2$ and $G^{(\infty)}\cong{\mathrm{SL}}_{2}(2^{f})$.
-In this case, the induced linear group of $G^{(\infty)}$ on $B_{f/2}$ is isomorphic to $\Omega^{-}_{4}(2^{f/2})$ (see [14, page 45]), which cannot be transitive on non-zero vectors of ${\mathbb{F}}_{2}^{2f}$.
+In this case, the induced linear group of $G^{(\infty)}$ on $B_{f/2}$ is isomorphic to $\Omega^{-}_{4}(2^{f/2})$ (see [16, page 45]), which cannot be transitive on non-zero vectors of ${\mathbb{F}}_{2}^{2f}$.
 Thus, $B_{j}\leqslant W$ for each $j$, and then $M$ is ${\mathbb{F}}_{2}G^{(\infty)}$ isomorphic to a quotient of $A\cong(\Lambda^{2}_{{\mathbb{F}}_{q}}(U))_{{\mathbb{F}}_{2}}$.
@@
   1. (i)
-    [9, Lemma 3] $K$ is an elementary abelian $p$-group of order $|{\mathbf{Z}}(N)|^{n}$;
+    [10, Lemma 3] $K$ is an elementary abelian $p$-group of order $|{\mathbf{Z}}(N)|^{n}$;
   2. (ii)
@@
 Thus, the number of $A$-orbits on $N$ is equal to $o(V)+o(M)-1$, as in part (iii).
-For $x,y,z\in N$, we notice that $[x,y]=[y,x]^{-1}$ and $[x,yz]=[x,y]^{z}[x,z]=[x,y][x,z]$ as $N^{\prime}={\mathbf{Z}}(N)$.
-Hence the commutator operator induces an alternative linear map:
+For $x,y,z\in N$, we notice that $[x,y]=[y,x]^{-1}$ and $[x,yz]=[x,y]^{z}[x,z]=[x,y][x,z]$ as $N^{\prime}={\mathbf{Z}}(N)$. Hence the commutator operator induces an alternating linear map:
 $\varphi:\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=V\wedge V\rightarrow M$ such that $\varphi:xC\wedge yC\mapsto[x,y]$.
@@
 Thus, ${\mathrm{Aut}}(N)$ is solvable, and $N$ is a Suzuki $2$-group by Shaw’s Theorem [23].
-The solvability of ${\mathrm{Aut}}(N)$ is proved by Bryukhanova in [3, Theorem 1.2].
+The solvability of ${\mathrm{Aut}}(N)$ is proved by Bryukhanova in [4, Theorem 1.2].
 ∎
 4. Finite $2$-groups with three fusion classes
-In this section, we introduce the three families of AT-$2$-groups, and then prove Corollary 1.3. The first family is $A_{2}(n,\theta)$, a Suzuki $2$-group of type A, see [11, page 82].
+In 2020, Bors and Glasby posed a paper [2] which classified finite $2$-groups with exactly three fusion classes.
+In this section, we introduce the families of $2$-groups with three fusion classes, and then obtain a similar classification as in [2] by using Theorem 1.1 and the main Theorem in [9].
+The first family is $A_{2}(n,\theta)$, a Suzuki $2$-group of type A, see [13, page 82].
 Definition 4.1.
@@
 $M(a,b,\theta)=\begin{pmatrix}1&a&b\\ 0&1&a^{\theta}\\ 0&0&1\end{pmatrix}.$
-We note that, if $n=2k+1$ and $x^{\theta}=x^{k+1}$, then $A_{2}(2k+1,\theta)$ is isomorphic to a Sylow $2$-subgroup of the Suzuki simple group ${\mathrm{Sz}}(2^{2k+1})$, see [6, Proposition 13.4(vi)].
-For any odd integers $n\neq 1$, the groups $A_{2}(n,\theta)$ are AT-groups by Zhang in [25, Lemma 2.2].
+We note that, if $n=2k+1$ and $x^{\theta}=x^{k+1}$, then $A_{2}(2k+1,\theta)$ is isomorphic to a Sylow $2$-subgroup of the Suzuki simple group ${\mathrm{Sz}}(2^{2k+1})$, see [7, Proposition 13.4(vi)].
+For any odd integers $n\neq 1$, the groups $A_{2}(n,\theta)$ are AT-groups by Zhang in [27, Lemma 2.2].
 The group $N=A_{2}(n,\theta)$ is a special $2$-group with ${\mathbf{Z}}(N)=\langle M(0,b,\theta)\mid b\in{\mathbb{F}}_{2^{n}}\rangle\cong{\mathbb{Z}}_{2}^{n}$.
@@
 ∎
-Sylow $2$-subgroups of ${\mathrm{SU}}_{3}(2^{n})$ are Suzuki $2$-group of type B (see [12, page 314]). Equip the space ${\mathbb{F}}_{q^{2}}^{3}$ ($q=2^{n}$) with the unitary form:
+Sylow $2$-subgroups of ${\mathrm{SU}}_{3}(2^{n})$ are Suzuki $2$-group of type B (see [14, page 314]). Equip the space ${\mathbb{F}}_{q^{2}}^{3}$ ($q=2^{n}$) with the unitary form:
 $\bigl{(}(x_{1},x_{2},x_{3}),(y_{1},y_{2},y_{3})\bigr{)}=x_{1}y_{3}^{q}+x_{2}y_{2}^{q}+x_{3}y_{1}^{q}.$
-Then the following group $B_{2}(n)$ is a Sylow $2$-subgroup of ${\mathrm{SU}}_{3}(q)$, see [7, page 249].
+Then the following group $B_{2}(n)$ is a Sylow $2$-subgroup of ${\mathrm{SU}}_{3}(q)$, see [8, page 249].
 Definition 4.3.
@@
 $M(a,b)=\begin{pmatrix}1&a&b\\ 0&1&a^{q}\\ 0&0&1\end{pmatrix}.$
-It was noticed in Zhang [25, Lemma 2.3] that $B_{2}(n)$ is an AT-group, with ${\mathbf{Z}}(N)=\langle M(0,b)\mid b\in{\mathbb{F}}_{q^{2}},b+b^{q}=0\rangle\cong{\mathbb{Z}}_{2}^{n}$.
+It was noticed in Zhang [27, Lemma 2.3] that $B_{2}(n)$ is an AT-group, with ${\mathbf{Z}}(N)=\langle M(0,b)\mid b\in{\mathbb{F}}_{q^{2}},b+b^{q}=0\rangle\cong{\mathbb{Z}}_{2}^{n}$.
 Lemma 4.4.
@@
 Let $D=\mathrm{diag}(\lambda^{-q},\lambda^{1-q},\lambda)$, where $\langle\lambda\rangle={\mathbb{F}}_{q^{2}}^{\times}$.
-Then $D^{-1}M(a,b)D=M(q\lambda,b\lambda^{q+1})$, and $b\lambda^{q+1}+(b\lambda^{q+1})^{q}+(a\lambda)(a\lambda)^{q}=\lambda^{1+q}(b+b^{q}+a^{1+q})=0$.
+Then $D^{-1}M(a,b)D=M(a\lambda,b\lambda^{q+1})$, and $b\lambda^{q+1}+(b\lambda^{q+1})^{q}+(a\lambda)(a\lambda)^{q}=\lambda^{1+q}(b+b^{q}+a^{1+q})=0$.
 Thus, $D$ induces an automorphism of $N$, denoted by $\xi$.
 Let $\phi$ be the Frobenius automorphism $x\mapsto x^{2}$ on ${\mathbb{F}}_{q^{2}}$, which induces an automorphism of $N$.
-As $M(a,b)$ and $M(q\lambda,b\lambda^{q+1})$ are not in the same coset of ${\mathbf{Z}}(N)$, it follows that $\langle\xi,\phi\rangle\cong{\Gamma\mathrm{L}}_{1}(2^{2n})$ acts faithfully on $N/{\mathbf{Z}}(N)$.
+As $M(a,b)$ and $M(a\lambda,b\lambda^{q+1})$ are not in the same coset of ${\mathbf{Z}}(N)$, it follows that $\langle\xi,\phi\rangle\cong{\Gamma\mathrm{L}}_{1}(2^{2n})$ acts faithfully on $N/{\mathbf{Z}}(N)$.
 Arguing as in the proof of Lemma 4.2, the kernel $K$ of ${\mathrm{Aut}}(N)$ on $N/{\mathbf{Z}}(N)$ is isomorphic to ${\mathbb{Z}}_{2}^{n^{2}}$, and ${\mathrm{Aut}}(N)/K=\langle\xi,\phi\rangle$.
@@
 Definition 4.5.
-For a generator $\epsilon$ of ${\mathbb{F}}_{2^{6}}^{\times}$, define
-$P(\epsilon)=\{(a,x)\in{\mathbb{F}}_{2^{6}}\times{\mathbb{F}}_{2^{3}}\mid(a,x)(b,y)=(a+b,x+y+ab^{2}\epsilon+a^{8}b^{16}\epsilon^{8})\}.$
+For a generator $\epsilon$ of ${\mathbb{F}}_{2^{6}}^{\times}$, define $P(\epsilon)$ to be the group consisting of elements in ${\mathbb{F}}_{2^{6}}\times{\mathbb{F}}_{2^{3}}$ with the opteration of multiplication defined by
+$(a,x)(b,w)=(a+b,x+y+ab^{2}\epsilon+a^{8}b^{16}\epsilon^{8}),\mbox{ for $(a,x),(b,w)\in{\mathbb{F}}_{2^{6}}\times{\mathbb{F}}_{2^{3}}$.}$
 The next lemma shows that the definition of $P(\epsilon)$ is independent of the choice of $\epsilon$.
@@
 Dornhoff [9] determined finite groups $N$ of which a solvable group of automorphisms has exactly three orbits on $N$.
-In his classification, desired non-abelian $2$-groups are precisely $A_{2}(n,\theta)$, $B_{p}(n)$ and $P(\epsilon)$.
-Then the following lemma immediately holds since each Suzuki $2$-group has a solvable automorphism group by Theorem 1.1(iii).
+In his classification, desired non-abelian $2$-groups are precisely $A_{2}(n,\theta)$, $B_{p}(n)$ and $P(\epsilon)$.= Then the following lemma immediately holds since each Suzuki $2$-group has a solvable automorphism group by Theorem 1.1(iii).
 Lemma 4.8.
 If $N$ is a finite non-abelian $2$-group such that ${\mathrm{Aut}}(N)$ has exactly three orbits on $N$, then $N=A_{2}(n,\theta)$, $B_{p}(n)$ or $P(\epsilon)$.
+The above lemma is the same as [2, Theorem 1.1] given by Bors and Glasby. They shows that $B_{p}(n)$ and $P(\epsilon)$ are in the family of Suzuki $2$-group of type B.
 Now we are ready to prove Corollary 1.3. Recall that two elements are conjugate under the automorphism group are said to be fused.